# Long-Tailed 3D Detection via Multi-Modal Late-Fusion

### v4

@@ Abstract
Contemporary autonomous vehicle (AV) benchmarks have advanced techniques for training 3D detectors, particularly on large-scale multi-modal (LiDAR + RGB) data. Surprisingly, although semantic class labels naturally follow a long-tailed distribution, existing benchmarks only focus on a few common classes (e.g., pedestrian and car) and neglect many rare but crucial classes (e.g., emergency vehicle and stroller). However, AVs must reliably detect both common and rare classes for safe operation in the open world. We address this challenge by formally studying the problem of Long-Tailed 3D Detection (LT3D), which evaluates all annotated classes, including those in-the-tail. We address LT3D with hierarchical losses that promote feature sharing across classes, and introduce diagnostic metrics that award partial credit to “reasonable” mistakes with respect to the semantic hierarchy (e.g., mistaking a child for an adult). Further, we point out that rare-class accuracy is particularly improved via multi-modal late fusion (MMLF) of independently trained uni-modal LiDAR and RGB detectors. Importantly, such an MMLF framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal detectors, unlike prevailing end-to-end trained multi-modal detectors that require paired multi-modal data. Finally, we examine three critical components of our simple MMLF approach from first principles and investigate whether to train 2D or 3D RGB detectors for fusion, whether to match RGB and LiDAR detections in 3D or the projected 2D image plane, and how to fuse matched detections. Extensive experiments reveal that 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors and matching on the 2D image plane mitigates depth estimation errors for better matching. Our proposed MMLF approach significantly improves LT3D performance over prior work, particularly improving rare class performance from 12.8 to 20.0 mAP! Our code and models are available on our project page.
Index Terms:
Long-Tailed Distribution, 3D Detection, Multi-Modal Late-Fusion, Autonomous Vehicles, Open World, LiDAR, RGB.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Figure 1: According to the histogram of per-class object counts (on the left), the nuScenes benchmark focuses on the common classes in cyan (e.g., car and barrier) but ignores rare ones in red (e.g., stroller and debris). In fact, the benchmark creates a superclass pedestrian by grouping multiple classes in green, including the common class adult and several rare classes (e.g., child and police-officer); this complicates the analysis of detection performance as pedestrian performance is dominated by adult. Moreover, the ignored superclass pushable-pullable also contains diverse objects such as shopping-cart, dolly, luggage and trash-can as shown in the top row (on the right). We argue that AVs should also detect rare classes as they can affect AV behaviors. Following [1], we report performance for three groups of classes based on their cardinality (split by dotted lines): Many, Medium, and Few.
Three-dimensional (3D) object detection is a key component in many robotics systems such as autonomous vehicles (AVs). To facilitate research in this space, the AV industry has released many large-scale 3D annotated multi-modal datasets [2, 3, 4, 5, 6]. However, these datasets often benchmark on only a few common classes (e.g., car and pedestrian) and ignore rare classes (although annotated by some datasets) like stroller and emergency vehicle (Fig. 1). In the real open world, safe navigation [7, 8] requires AVs to reliably detect rare-class objects such as debris and stroller. This motivates the study of Long-Tailed 3D Detection (LT3D), a problem that requires detecting objects from both common and rare classes.
@@
Status Quo. Among contemporary AV datasets, nuScenes [6] exhaustively annotates objects of various classes crucial to AVs (Fig. 1) and organizes them into a semantic hierarchy (Fig. 4). We are motivated to study LT3D by re-purposing all annotated classes in nuScenes because detecting rare classes is useful for downstream tasks such as motion planning. Importantly, LT3D is not simply solved by retraining state-of-the-art methods on both common and rare classes [9]. For example, CMT [10], a multi-modal transformer-based detector, achieves only 4.8 mAP on rare categories despite achieving 79.9 mAP on common classes (Table I).
@@
Protocol. LT3D requires 3D localization and recognition of objects from each of the common (e.g., adult and car) and rare classes (e.g, child and stroller). Moreover, for safety-critical robots such as AVs, detecting but mis-classifying rare objects (e.g., mis-classifying a child as an adult) is preferable to failing to detect them at all for safe operation. Therefore, we propose a new diagnostic metric to quantify the severity of classification mistakes in LT3D that exploits inter-class relationships w.r.t to the semantic hierarchy when awarding partial credit (Fig. 4). We use both the standard and proposed metrics to evaluate 3D detectors on all classes. Further, since prior work focuses on only a few common classes, they miss opportunities to exploit this semantic hierarchy during training. We propose hierarchical losses to promote feature sharing across both common and rare classes
@@
Figure 2: We extensively explore the simple multi-modal late-fusion (MMLF) framework for LT3D by ensembling RGB and LiDAR uni-modal detectors [9]. We rigorously examine three critical components within this framework (Fig. 3) and propose a simple method that fuses detections produced by a 2D RGB-detector (e.g., DINO [11]) and a 3D LiDAR-detector (e.g., CenterPoint [12]). Our method achieves 51.4 mAP on the nuScenes [6] LT3D benchmark, significantly improving over state-of-the-art detectors by 5.9% (Table I).
@@
Figure 3: We examine three key components in the multi-modal late-fusion (MMLF) of uni-modal RGB and LiDAR detectors from first principles: A. whether to train 2D or 3D RGB detectors, B. whether to match uni-modal detections on the 2D image plane or in the 3D bird’s-eye-view (BEV), and C. how to best fuse matched detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance (Table III and Table IV).
@@
Technical Insights. To address LT3D, we start by retraining state-of-the-art 3D detectors on all classes. Somewhat surprisingly, prior work performs rather poorly on rare classes, e.g., CenterPoint [12] achieves only 0.1 AP on child and 0.1 AP on stroller (Table V). We propose several algorithmic innovations to improve these results. First, we allow feature sharing across common and rare classes by training a single feature trunk, adding in hierarchical losses that ensure features will be useful for all classes (Table IV). Second, noting that LiDAR data is simply too impoverished for even humans to recognize certain tail objects that tend to be small (e.g., strollers), we propose a Multi-Modal Late-Fusion (MMLF) framework (Fig. 2) that fuses detections from a LiDAR-only detector (for precise 3D localization) and an RGB-only detector (for better recognition). We introduce Multi-Modal Filtering (MMF), a simple late-fusion approach that post-processes uni-modal 3D predictions from 3D LiDAR-only (e.g., CenterPoint [12]) and 3D RGB-only (e.g., FCOS3D [13]) detectors, filtering away detections that are inconsistent across modalities. This significantly improves LT3D performance by 3.2 mAP on average (Table II). Next, we delve into this MMLF framework and study three critical design choices (Fig. 3): (A) whether to train a 2D or a 3D monocular RGB detector for late-fusion, (B) whether to match detections in the 2D image plane or in 3D, and (C) how to optimally fuse detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining score-calibrated predictions with Bayesian fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal 3D detectors (Table I).
@@
Contributions. We present three major contributions. First, we formulate the problem of LT3D, emphasizing the detection of both common and rare classes in safety-critical applications like AVs. Second, we design LT3D’s benchmarking protocol and develop a diagnostic metric that awards partial credit depending on the severity of misclassifications (e.g., misclassifying child-vs-adult is less problematic than misclassifying child-vs-car). Third, we propose several architecture-agnostic approaches to address LT3D, including a simple multi-modal late-fusion (MMLF) strategy that generalizes across different RGB and LiDAR architectures. We conduct extensive experiments to ablate our design choices and demonstrate that our simple MMLF approach achieves state-of-the-art results on the nuScenes and Argoverse 2 LT3D benchmarks.
@@ 2 Related Work
3D Object Detection for AVs. Contemporary approaches to 3D object detection can be broadly classified as LiDAR-only, RGB-only, and sensor-fusion methods. Recent work in 3D detection is heavily inspired by prior work in 2D detection [14, 15, 16]. LiDAR-based detectors like PointPillars [17], CBGS [18], and PVRCNN++ [19] adopt an SSD-like architecture [15] that regresses amodal bounding boxes from a bird’s-eye-view (BEV) feature map. More recently, CenterPoint [12] adopts a center-regression loss that is inspired by CenterNet [14]. Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from the background given sparse LiDAR returns. Monocular RGB-based methods have gained popularity in recent years due to increased interest in camera-only perception. FCOS3D [13] extends FCOS [20] by additionally regressing the size, depth, and rotation for each object. More recently, methods such as BEVDet and BEVFormer [21, 22, 23] construct a BEV feature-map by estimating the per-pixel depth of each image feature [24]. PolarFormer [25] introduces a polar-coordinate transformation that improves near-field detection. Importantly, many of these state-of-the-art 3D RGB detectors are commonly pre-trained on large external datasets like DDAD [26]. Monocular RGB detectors accurately classify objects but struggle to estimate depth, particularly for far-field detections [27]. Despite recent advances in LiDAR and RGB 3D detectors, we find that multi-modal fusion is essential for LT3D (detailed next). Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect rare classes. We study the multi-modal late-fusion framework described in Fig. 2 to determine how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D.
@@
Multi-modal 3D Detection. Conventional wisdom suggests that fusing multi-modal cues, particularly using LiDAR and RGB, can improve 3D detection. Intuitively, LiDAR faithfully measures the 3D world (although it has notoriously sparse point returns), and RGB has high-resolution that captures semantic features for recognition (but lacks 3D information). Multi-modal fusion for 3D detection is an active field of exploration. Popular approaches can be categorized as input-fusion, feature-fusion, and late-fusion. Input-fusion methods typically augment LiDAR points using image-level features. For example, PointPainting [28] projects LiDAR points onto the output mask of a semantic segmentation model and appends corresponding class scores to each point. MVP [29] densifies regions of LiDAR sweeps that correspond with objects in semantic segmentation masks. In contrast, Frustum PointNets [30] leverage 2D RGB detections to localize objects within the box frustum using PointNets [31].
@@
Recent works show that feature-fusion can be more effective than input-fusion. PointFusion [32] fuses global image and point-cloud features prior to detection and MSMDFusion [33] fuses LiDAR and RGB features at multiple scales. TransFusion [34] and BEVFusion [35] fuse features in the BEV space using multi-headed attention. Despite the success of transformers for detecting common objects, [9] finds that TransFusion struggles to detect rare classes, and posits that the transformer architecture, as adopted in TransFusion and BEVFusion, suffers from limited training data (particularly for classes in the long tail). For transformers to work well in practice, they should be trained on diverse, large-scale datasets [36, 37]. It is also worth noting that end-to-end trained multi-modal detectors require paired multi-modal data for training, increasing the cost of data collection and modality alignment. In this work, we opt to study a multi-modal late-fusion strategy, a framework that ensembles uni-modal detectors, which do not require aligned RGB-LiDAR paired training data. CLOCs [38] is a late-fusion method that learns a separate network to fuse RGB and LiDAR detections, showing promising results for 3D detection. Importantly, prior late-fusion methods like CLOCS [38] only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications. We find that handling such misclassifications is critical for improving rare class performance. We delve into this simple multi-modal late-fusion framework, study three crucial design choices, and present a method that significantly outperforms the state-of-the-art for LT3D.
@@
Long-Tailed Perception. AV datasets follow a long-tailed class distribution: a few classes like car and pedestrian are dominant, while others like stroller and debris are rarely seen. However, this problem is not unique to the AV domain [39] – Long-Tailed Perception (LTP) is a long-standing problem in the literature [1] and has been widely studied through the lens of image classification, aiming for high accuracy averaged across imbalanced classes [1, 40, 41]. Existing LTP methods propose reweighting losses [42, 43, 44, 45, 46, 47], rebalancing data sampling [48, 49, 50], balancing gradients computed from imbalanced classes [51], and balancing network weights [41]. Others study LTP through the lens of 2D object detection with RGB images [52]. Compared to 2D image-based recognition, long-tailed 3D detection has unique opportunities and challenges because sensors such as LiDAR directly provide geometric and ego-motion cues that are difficult to extract from 2D images. Further, 2D detectors must detect objects of different scales due to perspective image projection, dramatically increasing the complexity of the output space (e.g., requiring more anchor boxes). In contrast, 3D objects do not exhibit as much scale variation, but far-away objects tend to have sparse LiDAR returns [27, 53], imposing different challenges. Finally, 3D detectors often use class-aware heads (i.e. each class has its own binary classifier) while 2D long-tail recognition approaches typically use shared softmax heads.
@@
Recently, CBGS [18] explicitly addresses rare-class 3D detection by up-sampling LiDAR-sweeps with instances of rare classes, and pasting instances of rare objects copied from different scenes. Although this works well for improving detection of infrequently-seen classes (e.g. classes with medium number of examples like bicycle and construction vehicle), it does not provide significant improvement for classes with only a few examples like debris and stroller [9]. Additionally, rare classes, such as child and stroller, are typically small in size and have a limited number of LiDAR returns. As a result, LiDAR-only detectors struggle to accurately recognize these rare classes. In this work, we address LT3D by multi-modal late-fusion, which ensembles RGB and LiDAR uni-modal detectors.

### v5

@@ Abstract
Contemporary autonomous vehicle (AV) benchmarks have advanced techniques for training 3D detectors, particularly on large-scale multi-modal (LiDAR + RGB) data. While class labels naturally follow a long-tailed distribution in the real world, surprisingly, existing benchmarks only focus on a few common classes (e.g., pedestrian and car) and neglect many rare but crucial classes (e.g., emergency vehicle and stroller). However, AVs must reliably detect both common and rare classes for safe operation in the open world. We address this challenge by formally studying the problem of Long-Tailed 3D Detection (LT3D), which evaluates all annotated classes, including those in-the-tail. We address LT3D with hierarchical losses that promote feature sharing across classes, and introduce diagnostic metrics that award partial credit to “reasonable” mistakes with respect to the semantic hierarchy (e.g., mistaking a child for an adult). Further, we point out that rare-class accuracy is particularly improved via multi-modal late fusion (MMLF) of independently trained uni-modal LiDAR and RGB detectors. Such an MMLF framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal detectors, unlike prevailing end-to-end trained multi-modal detectors that require paired multi-modal data. Finally, we examine three critical components of our simple MMLF approach from first principles: (1) whether to train 2D or 3D RGB detectors for fusion, (2) whether to match RGB and LiDAR detections in 3D or the projected 2D image plane, and (3) how to fuse matched detections. Extensive experiments reveal that (1) 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors, (2) matching on the 2D image plane mitigates depth estimation errors for better matching, and (3) score calibration and probabilistic fusion notably improves the final performance further. Our MMLF significantly outperforms prior work for LT3D, particularly improving on the six rarest classes from 12.8 to 20.0 mAP! Our code and models are available on our project page.
keywords:
Long-Tailed Distribution; 3D Detection; Multi-Modal Fusion; Late Fusion; Autonomous Vehicles; LiDAR; RGB.
@@ 1 Introduction
3D object detection is a key component in many robotics systems such as autonomous vehicles (AVs). To facilitate research in this space, the AV industry has released many large-scale 3D annotated multi-modal datasets [21, 7, 73, 64]. However, these datasets often benchmark on only a few common classes (e.g., car and pedestrian) and ignore rare classes (despite being annotated by some datasets [5]) like stroller and emergency vehicle (Fig. 1). In the real open world, safe navigation requires AVs to reliably detect rare classes such as debris and stroller [66, 75]. This motivates the study of Long-Tailed 3D Detection (LT3D), a problem that requires detecting objects from both common and rare classes.
@@
Figure 1: According to the histogram of per-class object counts (on the left), the nuScenes benchmark [5] focuses on the common classes in cyan (e.g., car and barrier) but ignores rare ones in red (e.g., stroller and debris). In fact, the benchmark creates a superclass pedestrian by grouping multiple classes in green, including the common class adult and several rare classes (e.g., child and police-officer); this complicates the analysis of detection performance as pedestrian performance is dominated by adult. Moreover, the ignored superclass pushable-pullable also contains diverse objects such as shopping-cart, dolly, luggage and trash-can as shown in the top row (on the right). We argue that AVs should also detect rare classes as they can affect AV behaviors. To comprehensively analyze performance, we report accuracy for three groups of classes (Many, Medium, and Few shown in the left panel) in the main paper and provide per-class performance in the appendix.
@@
Status Quo. Among contemporary AV datasets, nuScenes [5] exhaustively annotates objects of various classes crucial for safe navigation (Fig. 1) and organizes them into a semantic hierarchy (Fig. 4). We are motivated to study LT3D by re-purposing all annotated classes in nuScenes because detecting rare classes is useful for downstream tasks such as motion planning. Importantly, LT3D is not simply solved by retraining state-of-the-art methods on both common and rare classes [46]. For example, CMT [80], a multi-modal transformer-based detector, achieves only 4.8 mAP on rare categories despite achieving 79.9 mAP on common classes (Table 1). We significantly extend our previous conference paper [46] which introduces LT3D by (1) studying whether to fuse 2D or 3D RGB detectors; (2) evaluating whether to fuse in 2D image plane or 3D LiDAR space; (3) investigating the impact of score calibration and probabilistic fusion; (4) rigorously comparing recently published end-to-end trained multimodal detectors.
@@
Protocol. LT3D requires 3D localization and recognition of objects from each of the common (e.g., adult and car) and rare classes (e.g, child and stroller). Moreover, for safety-critical robots such as AVs, detecting but mis-classifying rare objects (e.g., mis-classifying a child as an adult) is preferable to failing to detect them at all. Therefore, we propose a new diagnostic metric to quantify the severity of classification mistakes in LT3D that exploits inter-class relationships w.r.t to the semantic hierarchy when awarding partial credit (Fig. 4). We use both the standard and proposed metrics to evaluate 3D detectors on all classes. Further, as prior works focus on only a few common classes, they miss opportunities to exploit this hierarchy in training. We propose hierarchical losses to promote feature sharing across common and rare classes, achieving improved performance.
@@
Figure 2: We extensively explore the simple multi-modal late-fusion (MMLF) framework for LT3D by fusing RGB and LiDAR uni-modal detectors [46]. We rigorously examine three critical components within this framework (Fig. 3) and propose a simple method that fuses detections produced by a 2D RGB-detector (e.g., DINO [88]) and a 3D LiDAR-detector (e.g., CenterPoint [82]). Our method achieves 51.4 mAP on the nuScenes [5] LT3D benchmark, significantly improving over state-of-the-art detectors by 5.9% (Table 1).
@@
Figure 3: We examine three key components in our multi-modal late-fusion (MMLF) of uni-modal RGB and LiDAR detectors from first principles: A. whether to train 2D or 3D RGB detectors, B. whether to match uni-modal detections on the 2D image plane or in the 3D bird’s-eye-view (BEV), and C. how to best fuse matched detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance (Table 3 and Table 4).
@@
Technical Insights. To address LT3D, we start by retraining state-of-the-art 3D detectors on all classes. Somewhat surprisingly, these detectors perform rather poorly on rare classes, e.g., CenterPoint [82] achieves only 0.1 AP on child and stroller (Table 5). We propose several algorithmic innovations to improve these results. First, we allow feature sharing across common and rare classes by training a single feature trunk, adding in hierarchical losses that ensure features will be useful for all classes (Table 4). Second, noting that LiDAR data is too impoverished for even humans to recognize certain rare objects that tend to be small (e.g., strollers), we propose a Multi-Modal Late-Fusion (MMLF) framework (Fig. 2) that fuses detections from a LiDAR-only detector (for precise 3D localization) and an RGB-only detector (for better recognition). Within MMLF, we first introduce a simple Multi-Modal Filtering (MMF) method, which post-processes predictions from 3D LiDAR-only (e.g., CenterPoint [82]) and 3D RGB-only (e.g., FCOS3D [76]) detectors, filtering away detections that are inconsistent across modalities. This improves LT3D performance by 3.2 mAP (Table 2). Next, we delve into this MMLF framework and study three critical design choices (Fig. 3): (A) whether to train a 2D or a 3D RGB detector for late-fusion, (B) whether to match detections in the 2D image plane or in 3D, and (C) how to optimally fuse detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining score-calibrated predictions with probabilistic fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal 3D detectors (Table 1).
@@
Contributions. We present three major contributions. First, we formulate the problem of LT3D, emphasizing the detection of both common and rare classes in safety-critical applications like AVs. Second, we design LT3D’s benchmarking protocol and develop a diagnostic metric that awards partial credit depending on the severity of misclassifications (e.g., misclassifying child-vs-adult is less problematic than misclassifying child-vs-car). Third, we propose several architecture-agnostic approaches to address LT3D, including a simple multi-modal late-fusion (MMLF) strategy that generalizes across different RGB and LiDAR architectures (Appendix D). We conduct extensive experiments to ablate our design choices and demonstrate that our simple MMLF approach achieves state-of-the-art results on the nuScenes and Argoverse 2 LT3D benchmarks.
@@ 2 Related Work
3D Object Detection for AVs. Contemporary approaches to 3D object detection can be broadly classified as LiDAR-only, RGB-only, and sensor-fusion methods. Recent work in 3D detection is inspired by prior work in 2D detection [86, 33, 8]. LiDAR-based detectors like PointPillars [41], CBGS [84], and PVRCNN++ [63] adopt an SSD-like architecture [33] that regresses amodal bounding boxes from a bird’s-eye-view (BEV) feature map. More recently, CenterPoint [82] adopts a center-regression loss that is inspired by CenterNet [86]. Recent LiDAR-based detectors adopt the transformer architecture or attention modules that achieve further improvements [30, 34]. Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from the background given sparse LiDAR returns. Monocular RGB-based methods have gained popularity in recent years due to increased interest in camera-only perception. FCOS3D [76] extends FCOS [67] by additionally regressing the size, depth, and rotation for each object. More recently, methods such as BEVDet and BEVFormer [25, 24, 43] construct a BEV feature-map by estimating the per-pixel depth of each image feature [47]. PolarFormer [31] introduces a polar-coordinate transformation that improves near-field detection. Many of these state-of-the-art 3D RGB detectors are commonly pre-trained on large external datasets like DDAD [17]. Monocular RGB detectors accurately classify objects but struggle to estimate depth, particularly for far-field detections [20]. Despite recent advances in LiDAR and RGB 3D detectors, we find that multi-modal fusion is essential for LT3D (detailed next). Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect rare classes. Our work focuses on multi-modal late-fusion (Fig. 2) and studies how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D.
@@
Multi-modal 3D Detection. Multi-modal fusion for 3D detection is an active field of exploration [2, 40, 81]. Existing methods can be categorized as input-fusion, feature-fusion, and late-fusion. Input-fusion typically augments LiDAR points using image-level features. For example, PointPainting [69] projects LiDAR points onto the output mask of a semantic segmentation model and appends corresponding class scores to each point. MVP [83] densifies regions of LiDAR sweeps that correspond with objects in semantic segmentation masks. In contrast, Frustum PointNets [50] leverage 2D RGB detections to localize objects within the box frustum using PointNets [51].
@@
Recent works show that feature-fusion is more effective than input-fusion. [77] fuses global image and point-cloud features prior to detection; [29, 11] fuse LiDAR and RGB features at multiple scales. TransFusion [2] and BEVFusion [40] fuse features in the BEV space using multi-headed attention. Despite the success of transformers for detecting common objects, [46] finds that TransFusion struggles to detect rare classes, and posits that the transformer architecture, as adopted in TransFusion and BEVFusion, suffers from limited training data (particularly for classes in the long tail). For transformers to work well in practice, they should be trained on diverse, large-scale datasets [12, 59]. It is also worth noting that end-to-end trained multi-modal detectors require paired multi-modal data for training, increasing the cost of data collection and modality alignment. In contrast, our work focuses on multi-modal late-fusion (MMLF), a framework that ensembles uni-modal detectors, which do not require aligned RGB-LiDAR paired training data. CLOCs [49] is a late-fusion method that learns a separate network to fuse RGB and LiDAR detections, showing promising results for 3D detection. However, CLOCS only performs late fusion on matched predictions with semantic agreement and does not fix misclassifications. We find that handling such misclassifications is critical for improving rare class performance. We delve into this simple MMLF framework, study three crucial design choices, and present the final method that significantly outperforms the state-of-the-art approaches for LT3D.
@@
Long-Tailed Perception. AV datasets follow a long-tailed class distribution: a few classes like car and pedestrian are dominant, while others like stroller and debris are rarely seen. In fact, Long-Tailed Perception (LTP) is not a unique problem in the AV domain [54] but a long-standing problem in the literature [37] and has been widely studied through the lens of image classification, aiming for high accuracy averaged across imbalanced classes [37, 85, 1]. Existing LTP methods propose reweighting losses [6, 32, 26, 89], rebalancing data sampling [14, 4, 27], balancing gradients computed from imbalanced classes [65], and balancing network weights [1]. Others study LTP through the lens of 2D object detection [19]. Compared to long-tailed image classification, long-tailed 3D detection provides unique opportunities and challenges as sensors such as LiDAR directly provide geometric and ego-motion cues that are difficult to extract from 2D images. Further, 2D detectors must detect objects of different scales due to perspective image projection, dramatically increasing the complexity of the output space (e.g., requiring more anchor boxes). In contrast, 3D objects do not exhibit as much scale variation, but far-away objects tend to have sparse LiDAR returns [20, 48], imposing different challenges. Finally, 3D detectors often use class-aware heads (i.e. each class has its own binary classifier) while 2D long-tail recognition approaches typically use shared softmax heads.
@@
Recently, CBGS [84] explicitly addresses rare-class 3D detection by up-sampling LiDAR-sweeps with instances of rare classes and pasting instances of rare objects copied from different scenes. Although this works well for improving detection of infrequently-seen classes (e.g. classes with medium number of examples like bicycle and construction vehicle), it does not provide significant improvement for classes with only a few examples like debris and stroller [46]. Additionally, rare classes, such as child and stroller, are typically small in size and have a limited number of LiDAR returns. As a result, LiDAR-only detectors struggle to accurately recognize these rare classes. In this work, we address LT3D by multi-modal late-fusion, which ensembles RGB and LiDAR uni-modal detectors.